OpenStudy (anonymous):

WILL GIVE MEDAL IF YOU ANSWER THESE QUESTIONS CORRECTLY! For the function y = f(x), what is the ordered pair for the point on the graph when x = 2m + 3? (3 points) (x, f(m)) (x, 2m + 3) (2m + 3, f(m)) (2m + 3, f(2m + 3)),

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

by any chance do you go to a school by the name of David Posnack

OpenStudy (solomonzelman):

the ordered pair is (x,y) (same as (x,f(x) ) but in this case, you are given that x is 2m+3, so the ordered pair would be, \(\large\color{black}{(2m + 3, f(2m + 3)) }\)

OpenStudy (solomonzelman):

No, I go to college. (Some college in Atlanta Georgia)

OpenStudy (anonymous):

ok thanks can you answer a few more?

OpenStudy (anonymous):

The reason why i ask is because your name is so familiar

OpenStudy (anonymous):

Given the function g(x) = -4x + 5, compare and contrast g(2) and g(-4). Choose the statement that is true concerning these two values. (3 points) The value of g(2) is larger than the value of g(-4). The value of g(2) is smaller than the value of g(-4). The value of g(2) is the same as the value of g(-4). The values of g(2) and g(-4) cannot be compared.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

We will do this together... \(\large\color{black}{ g(x) = -4x + 5}\) When you say, \(\large\color{black}{ g(2)}\) you are finding the function when x=2, so plug in 3 for x. Same way, plug in -4 to find \(\large\color{black}{ g(-4) }\) \(\large\color{black}{ g(-4) = -4(-4) + 5}\) \(\large\color{black}{ g(-4)= 16 + 5}\) \(\large\color{black}{ g(-4)= 21}\) \(\large\color{black}{ g(2)= -4(2) + 5}\) \(\large\color{black}{ g(2)= -8 + 5}\) \(\large\color{black}{ g(2)= -3}\) So which one is larger ?

OpenStudy (anonymous):

g4?

OpenStudy (anonymous):

sorry (g-4)

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

g(2) is smaller, so the answer is B.

OpenStudy (solomonzelman):

yes ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1 more

OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? (4 points) -2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180, …

OpenStudy (anonymous):

andWhich of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) + 9? (3 points) The graph of y = f(x) will shift up 9 units. The graph of y = f(x) will shift down 9 units. The graph of y = f(x) will shift left 9 units. The graph of y = f(x) will shift right 9 units.

OpenStudy (anonymous):

?